any recent questions and answers from the tease about mathematics and science
Answer #1

ATI teas science questions :-

- Alveoli are tiny sacs surrounded by what?

capillaries

- Veins are blood vessels that transport blood from the capillaries to the what?

heart

- Arteries are blood vessels that transport blood away from the heart and to the what?

capillaries

- What serves as the location for the exchange of oxygen, carbon dioxide fluid and nutrients within the body?

capillaries

- What are the bronchial tubes are lined with

cilia

- What circulatory system includes what body structures?

heart, blood, blood vessels

- Deoxygenated blood enters the heart through which two structures?

1) Superior vena cava

2) Inferior vena cava

- When the right atrium contracts, blood flows through the tricuspid valve and goes where?

right ventricle

- How many chambers does the heart have & what are the names of the chambers?

The heart has 4 chambers.

1-Right Atrium

2-left Atrium

3-Right ventricle

4-Left ventricle

- Inhaled air travels from the trachea into what structure?

the right and left bronchial tubes

#. Maths questions :-

- If Lynn can type a page in p minutes, what piece of the page can she do in 5 minutes?

5/p

- If Sally can paint a house in 4 hours, and John can paint the same house in 6 hour, how long will it take for both of them to paint the house together?

2 hours and 24 minutes

- Employees of a discount appliance store receive an additional 20% off of the lowest price on an item. If an employee purchases a dishwasher during a 15% off sale, how much will he pay if the dishwasher originally cost $450?

Sale Price = $450 - 0.15$450 = $382.50, Employee Price = $382.50 - 0.2$382.50 = $306

- The sales price of a car is $12,590, which is 20% off the original price. What is the original price?

$12,590 = Original Price - 0.2Original Price = 0.8Original Price, Original Price = $12,590/0.8 = $15,737.50

- Solve the following equation for A : 2A/3 = 8 + 4A

-2.4

- If Leah is 6 years older than Sue, and John is 5 years older than Leah, and the total of their ages is 41. Then how old is Sue?

8

- Alfred wants to invest $4,000 at 6% simple interest rate for 5 years. How much interest will he receive?

$1,200

- Jim is able to sell a hand-carved statue for $670 which was a 35% profit over his cost. How much did the statue originally cost him?

$670 = Cost + 0.35Cost = 1.35Cost, Cost = $670/1.35 = $496.30

- The city council has decided to add a 0.3% tax on motel and hotel rooms. If a traveler spends the night in a motel room that costs $55 before taxes, how much will the city receive in taxes from him?

17 cents

- A student receives his grade report from a local community college, but the GPA is smudged. He took the following classes: a 2 hour credit art, a 3 hour credit history, a 4 hour credit science course, a 3 hour credit mathematics course, and a 1 hour science lab. He received a "B" in the art class, an "A" in the history class, a "C" in the science class, a "B" in the mathematics class, and an "A" in the science lab. What was his GPA if the letter grades are based on a 4 point scale?

3.0